# Deflated Dynamics Value Iteration

### v1

@@ 3.2.  Matrix Splitting: SOR
$z=(D+\alpha B)^{-1}(\alpha b-(\alpha C-(\alpha-1)D)z),$
provided that $D+\alpha B$ is invertible. Successive over-relaxation (SOR) attempts to find a solution through the fixed-point iteration
$z^{k+1}=(D+\alpha B)^{-1}(\alpha b-(\alpha C-(\alpha-1)D)z^{k}).$
@@
Policy evaluation is the problem of finding a unique $V^{\pi}$ that satisfies the Bellman equation $V^{\pi}=T^{\pi}V^{\pi}$, or in an expanded form, $V^{\pi}=r^{\pi}+\gamma\mathcal{P}^{\pi}V^{\pi}$. Notice that this is equivalent to
$(I-\gamma E-\gamma(\mathcal{P}^{\pi}-E))V^{\pi}=r^{\pi}$
for any $E\in\mathbb{R}^{n\times n}$. The SOR iteration for PE is
$V^{k+1}=(I-\alpha\gamma E)^{-1}(\alpha r^{\pi}+((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E))V^{k}).$
As an example, notice that for $\alpha=1$ and $E=0$, we recover the original VI. The convergence behaviour of this procedure depends on the choice of $\alpha$ and $E$. Soon we propose particular choices that can lead to acceleration.
3.3.  Deflated Dynamics Value Iteration
We are now ready to introduce DDVI. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$ satisfying the conditions of Facts 1, 2, or 3. Using (1), we can express the SOR iteration for PE with deflation matrix $E_{s}$ as
@@ 3.3.  Deflated Dynamics Value Iteration
Let $\pi$ be a policy and let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. Let $s\leq n$. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$ satisfying the conditions of Facts 1, 2, or 3. For $0<\alpha\leq 1$, DDVI (2) exhibits the rate [^1] The precise meaning of the $\tilde{O}$ notation is $\|V^{k}-V^{\pi}\|=O(|\lambda+\epsilon|^{k}\|V^{0}-V^{\pi}\|)$ as $k\rightarrow\infty$ for any $\epsilon>0$.
@@
When $\alpha=1$, we can simplify DDVI (2) as follows.
@@
In the setting of Theorem 3.1, if $\alpha=1$, then DDVI (2) simplifies to
$W^{k+1}=\gamma(\mathcal{P}^{\pi}-E_{s})V^{k}+r^{\pi},\quad V^{k+1}=\left(I+\sum^{s}_{i=1}\frac{\gamma\lambda_{i}}{1-\gamma\lambda_{i}}u_{i}v_{i}^{\top}\right)W^{k+1}$
and the the rate reduces to $\|V^{k}-V^{\pi}\|=\tilde{O}(|\gamma\lambda_{s+1}|^{k}\|V^{0}-V^{\pi}\|)$ as $k\rightarrow\infty$.
Let us compare this convergence rate with the original VI’s. The rate for VI is $O(\gamma^{k})$, which is slow when $\gamma\approx 1$. By choosing $E$ to deflate the top $s$ eigenvalues of $\mathcal{P}^{\pi}$, DDVI has the rate of $O(|\gamma\lambda_{s+1}|^{k})$ behaviour, which is exponentially faster whenever $\lambda_{s+1}$ is smaller than $1$. The exact behaviour depends on the spectrum of the Markov chain (whether eigenvalues are close to $1$ or far from them) and the number of eigenvalues we decided to deflate by $E$.
In Section 4, we discuss a practical implementation of DDVI based on the power iteration and orthogonal iteration and implement it in the experiments of Section 6. We find that smaller values of $\alpha$ lead to more stable iterations when using approximate deflation matrices. We also show how we can use DDVI in the RL setting in Section 5.
This version of DDVI (2) applies to the policy evaluation (PE) setup. The challenge in extending DDVI to the Control setup is that $\mathcal{P}^{\pi}$ changes throughout the iterations, and so should the deflation matrix $E_{s}$. However, when $s=1$, the deflation matrix $E_{1}$ can be kept constant, and we utilize this fact in Section 4.1.
4.  Computing Deflation Matrix $E$
@@
4.1.  Rank-$1$ DDVI for PE and Control
@@
Here, we benefitted from the fact that $E_{1}$ is not a function of $\pi$ in order to take it out of the $\max_{\pi}$. Compare with (2) of DDVI (PE), we set $\alpha=1$ for simplicity. We have the following guarantee for rank-$1$ DDVI for Control.
@@
The DDVI for Control algorithm (3) exhibits the rate $\|V^{k}-V^{\star}\|\leq\frac{2}{1-\gamma}\gamma^{k}\|V^{0}-V^{\star}\|$, for $k=0,1,\dots$ where $V^{k}=(I+\frac{\gamma}{1-\gamma}\mathbf{1}v^{\top})W^{k}$. Furthermore, if there exist a unique optimal policy $\pi^{\star}$, then $\|V^{k}-V^{\star}\|=\tilde{O}(|\gamma\lambda_{2}|^{k}\|V^{0}-V^{\star}\|)$ as $k\rightarrow\infty$, where $\lambda_{2}$ is the second eigenvalue of $\mathcal{P}^{\pi^{\star}}$ (The $\tilde{O}$ notation is as defined in Theorem 3.1).
@@ Discussion.
Although rank-1 DDVI for Control does accelerate the convergence $V^{k}\rightarrow V^{\star}$, a subtle point to note is that the greedy policy $\pi^{k}$ obtained from $V^{k}$ is not affected by the rank-1 deflation. Briefly speaking, this is because adding a uniform constant to $V^{k}$ through $\mathbf{1}$ has no effect on the $\operatorname*{argmax}$. Indeed, the maximizer $\pi^{k+1}$ in (3) is the same as $\operatorname*{argmax}_{\pi}\{r^{\pi}+\gamma{\mathcal{P}}^{\pi}V^{k}\}$, produced by the (non-deflated) value iteration, when $W^{0}=V^{0}$. Having the term $v^{\top}\mathbf{1}W^{k}$ in the update $W^{k+1}=r^{\pi}+\gamma(\mathcal{P}^{\pi^{k+1}}-v^{\top}\mathbf{1})W^{k}$ adds the same constant to all states, so it does not change the maximizer of the next policy.
4.2.  DDVI with Automatic Power Iteration
@@
Assume that the top $s+1$ eigenvalues of $\mathcal{P}^{\pi}$ have distinct magnitude, i.e., $1=\lambda_{1}>|\lambda_{2}|>\cdots>|\lambda_{s+1}|$. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$ as in Fact 2.
@@
Consider the DDVI algorithm (2) with $\alpha=1$ as in Corollary 3.2. If $V^{0}=0$, since $\left(\mathcal{P}^{\pi}-E_{s}\right)(I-\gamma E_{s})^{-1}=\mathcal{P}^{\pi}-E_{s}$ (verified in Appendix 8), we have $W^{k}=\sum_{i=0}^{k-1}\gamma^{i}\left(\mathcal{P}^{\pi}-E_{s}\right)^{i}r^{\pi}$. Therefore, $\frac{1}{\gamma^{k}}(W^{k+1}-W^{k})=\left(\mathcal{P}^{\pi}-E_{s}\right)^{k}r^{\pi}$ is the iterates of a power iteration for the matrix $(\mathcal{P}^{\pi}-E_{s})$ starting from initial vector $r^{\pi}$. As discussed in Fact 2, the top eigenvalue of $(\mathcal{P}^{\pi}-E_{s})$ is $\lambda_{s+1}$. For large $k$, we expect $\frac{W^{k+1}-W^{k}}{\|W^{k+1}-W^{k}\|}\approx w$, where $w$ is the top eigenvector of $(\mathcal{P}^{\pi}-E_{s})$. With $w$, we can recover the $(s+1)$-th right eigenvector of $\mathcal{P}^{\pi}$ through the formula ([Bru et al., 2012, Proposition 5]) $u_{s+1}=w-\sum^{s}_{i=1}\frac{\lambda_{i}v_{i}^{\top}w}{\lambda_{i}-\lambda_{s+1}}u_{i}$.
@@
Leveraging this observation, DDVI with Automatic Power Iteration (AutoPI) computes an approximate rank-$s$ deflation matrix $E_{s}$ while performing DDVI: Start with a rank-$1$ deflation matrix, using the first right eigenvector $u_{1}=\mathbf{1}$, and carry out DDVI iterations. If a certain error criteria is satisfied, use $W^{k+1}-W^{k}$ to approximate the second right eigenvector $u_{2}$. Then, update the deflation matrix to be rank $2$, and gradually increase the deflation rank. Algorithm 3 in Appendix 9 has the detail.
Rank-$s$ DDVI with the QR Iteration.
We can also use the QR Iteration to compute the top $s$ Schur vectors, and then construct rank-$s$ Schur deflation matrix and performs DDVI. We can also perform the QR Iteration in an automated manner, similar to AutoPI. This results in the AutoQR algorithm. The detail is in Appendix 9.
5.  Deflated Dynamics Temporal Difference Learning
@@
Many practical RL algorithms such as Temporal Difference (TD) Learning [Sutton, 1988, Tsitsiklis and V. Roy, 1997], Q-Learning [Watkins, 1989], Fitted Value Iteration [Gordon, 1995], and DQN [Mnih et al., 2015] can be viewed as sample-based variants of VI. Consequently, the slow convergence in the case of $\gamma\approx 1$ has also been observed for these algorithms by Szepesvári [1997], Even-Dar and Mansour [2003], Wainwright [2019] for TD Learning and by Munos and Szepesvári [2008], Farahmand et al. [2010], Chen and Jiang [2019], Fan et al. [2020] for Fitted VI and DQN. Here we introduce Deflated Dynamics Temporal Difference Learning (DDTD) as a sample-based variant of DDVI. To start, recall that the tabular temporal difference (TD) learning performs the updates
$V^{k+1}(X_{k})=V^{k}(X_{k})+\eta_{k}[r^{\pi}(X_{k})+\gamma V^{k}(X^{\prime}_{k})-V^{k}(X_{k})],$
@@
where $\eta_{k}$ is stepsize, $(X_{k},X^{\prime}_{k})$ are random samples from the environment such that $X_{k}^{\prime}$ is the subsequent state following $X_{k}$, and $r^{\pi}(X_{k})$ is the expected $1$-step reward obtained by following policy $\pi$ from state $X_{k}$. TD is a sample-based variant of VI for PE. Two key ingredients of TD learning are the random coordinate updates and the temporal difference error $r^{\pi}(X)+\gamma V(X^{\prime})-V(X)$ whose conditional expectation is equal to $(T^{\pi}V-V)(X)$. Applying the same ingredients to DDVI, we obtain DDTD. The improved convergence rate for DDVI compared to VI suggests that DDTD will also exhibit improved convergence rates.
@@
$W^{k+1}(X_{k})$ $=W^{k}(X_{k})+\eta_{k}\big{[}(1-\alpha)V^{k}(X_{k})+\alpha\left(r^{\pi}(X_{k})+\gamma V(X^{\prime}_{k})-\gamma(E_{s}V^{k})(X_{k})\right)-W^{k}(X_{k})\big{]}$
@@
Let $\eta_{k}=(\sum^{k}_{i=0}\mathbf{1}_{X_{i}=x})^{-1}$ when $X_{k}=x$. For $\alpha=1$, DDTD (4) converges to $V^{\pi}$ almost surely.
@@
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. Let $\eta_{k}=C/(k+1)$ for $C>\frac{n}{2\lambda_{\text{DDTD}}}$ where $\lambda_{\text{DDTD}}=\min_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\text{Re}(1-\gamma\lambda)$. For $\alpha=1$, DDTD (4) exhibits the rate $\mathbf{E}[\|V^{k}-V^{\pi}\|^{2}]=O(k^{-1})$.
@@
5.1.  Implementation of DDTD
To implement DDTD practically, we take a model-based approach for obtaining $E_{s}$. At each iteration, the agent uses the new samples to update an approximate model $\hat{\mathcal{P}}^{\pi}$ of the true dynamics. This updated approximate model is used to compute $E_{s}$. We update $E_{s}$ every $K$-th iterations since the change in $\hat{\mathcal{P}}^{\pi}$ and consequently the resulting $E_{s}$ at each iteration is small. Whenever a new $E_{s}$ is computed, we set $W$ to be $(I-\alpha\gamma E_{s})V$. This step ensures that $V$ smoothly converges to $V^{\pi}$ by updating all coordinates of $W$ at once based on new $E_{s}$. Also, we use the random sample reward $R$ in place of the expected reward $r^{\pi}(X_{k})$. We formally state the algorithm as Algorithm 1.

### v2

@@ 3.2 Matrix Splitting: Successive over-relaxation
$z=(D+\alpha B)^{-1}(\alpha b-(\alpha C+(\alpha-1)D)z),$
provided that $D+\alpha B$ is invertible. SOR attempts to find a solution through the fixed-point iteration
$z^{k+1}=(D+\alpha B)^{-1}(\alpha b-(\alpha C+(\alpha-1)D)z^{k}).$
@@
In the case of the policy evaluation problem of equation 2, we have $A=I-\gamma\mathcal{P}^{\pi}$ and $b=r^{\pi}$. For any $E\in\mathbb{R}^{n\times n}$, we can consider the splitting
$I-\gamma\mathcal{P}^{\pi}=-\gamma E-\gamma(\mathcal{P}^{\pi}-E)+I,$
where we chose $B=-\gamma E$, $C=-\gamma(\mathcal{P}^{\pi}-E)$, and $D=I$. With these choices, the SOR iteration for PE is
$V^{k+1}=(I-\alpha\gamma E)^{-1}(\alpha r^{\pi}+((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E))V^{k}).$ (3)
Notice that $V^{\pi}$ is the fixed point of this iterative procedure: $V^{\pi}=(I-\alpha\gamma E)^{-1}(\alpha r^{\pi}+((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E))V^{\pi})$. Whether these iterations converge to the fixed point, however, depends on the choice of $\alpha$ and $E$. As an example, for $\alpha=1$ and $E=0$, we recover the original VI, which is convergent with the convergence rate of $O(\gamma^{k})$. Next, we propose particular choices that lead to not only convergence but acceleration.
3.3 Deflated Dynamics Value Iteration
We are now ready to introduce DDVI. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$ satisfying one of the conditions of Facts 1, 2, or 3. The SOR iteration for PE with deflation matrix $E_{s}$ (cf. equation 3) is
$V^{k+1}=(I-\alpha\gamma E_{s})^{-1}(\alpha r^{\pi}+((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))V^{k}).$ (4)
By benefitting from equation 1, we can write it as
@@ 3.3 Deflated Dynamics Value Iteration
Let $\pi$ be a policy and let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. Let $s\leq n$. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$ satisfying the conditions of Facts 1, 2, or 3. For $0<\alpha\leq 1$, DDVI equation 5 exhibits the rate [^1] The precise meaning of the $\tilde{O}$ notation is $\|V^{k}-V^{\pi}\|=O(|\lambda+\epsilon|^{k}\|V^{0}-V^{\pi}\|)$ as $k\rightarrow\infty$ for any $\epsilon>0$.
@@
When $\alpha=1$, we can simplify DDVI equation 5 as follows.
@@
In the setting of Theorem 3.1, if $\alpha=1$, then DDVI equation 5 simplifies to
$W^{k+1}=\gamma(\mathcal{P}^{\pi}-E_{s})V^{k}+r^{\pi},$
$V^{k+1}=\left(I+\sum^{s}_{i=1}\frac{\gamma\lambda_{i}}{1-\gamma\lambda_{i}}u_{i}v_{i}^{\top}\right)W^{k+1},$
and the the rate simplifies to $\|V^{k}-V^{\pi}\|=\tilde{O}(|\gamma\lambda_{s+1}|^{k}\|V^{0}-V^{\pi}\|)$ as $k\rightarrow\infty$.
Let us compare this convergence rate with the original VI’s. The rate for VI is $O(\gamma^{k})$, which is slow when $\gamma\approx 1$. By choosing $E_{s}$ to deflate the top $s$ eigenvalues of $\mathcal{P}^{\pi}$, DDVI has the rate of $O(|\gamma\lambda_{s+1}|^{k})$, which is exponentially faster whenever $\lambda_{s+1}$ is smaller than $1$. The exact behaviour depends on the spectrum of the Markov chain (whether eigenvalues are close to $1$ or far from them) and the number of eigenvalues we decided to deflate by $E_{s}$.
In Section 4, we discuss a practical implementation of DDVI based on the power iteration and QR iteration and implement it in the experiments of Section 6. We find that smaller values of $\alpha$ lead to more stable iterations when using approximate deflation matrices. We also show how we can use DDVI in the RL setting in Section 5.
This version of DDVI equation 5 applies to the policy evaluation (PE) setup. The challenge in extending DDVI to the Control setup is that $\mathcal{P}^{\pi}$ changes throughout the iterations, and so should the deflation matrix $E_{s}$. However, when $s=1$, the deflation matrix $E_{1}$ can be kept constant, and we utilize this fact in Section 4.1.
4 Computing Deflation Matrix $E$
@@
4.1 Rank-$1$ DDVI for PE and Control
@@
Here, we benefitted from the fact that $E_{1}$ is not a function of $\pi$ in order to take it out of the $\max_{\pi}$. The value function $V$ can be computed as
$V^{k}=\left(I+\frac{\gamma}{1-\gamma}\mathbf{1}v^{\top}\right)W^{k}.$ (7)
Note that the iteration over $W^{k}$ does not depend on $V^{k}$, so we do not need to compute $V$ at each iteration – only when we need it. Compared to equation 5 of DDVI (PE), we set $\alpha=1$ for simplicity. We have the following guarantee for rank-$1$ DDVI for Control.
@@
The DDVI for Control algorithm (equation 6 and equation 7) exhibits the rate
$\|V^{k}-V^{\star}\|\leq\frac{2}{1-\gamma}\gamma^{k}\|V^{0}-V^{\star}\|,$
for $k=0,1,\dots$. Furthermore, if there exist a unique optimal policy $\pi^{\star}$, then
$\|V^{k}-V^{\star}\|=\tilde{O}(|\gamma\lambda_{2}|^{k}\|V^{0}-V^{\star}\|)$
as $k\rightarrow\infty$, where $\lambda_{2}$ is the second eigenvalue of $\mathcal{P}^{\pi^{\star}}$ (The $\tilde{O}$ notation is as defined in Theorem 3.1).
@@ Discussion.
Although rank-1 DDVI for Control does accelerate the convergence $V^{k}\rightarrow V^{\star}$, a subtle point to note is that the greedy policy $\pi^{k}$ obtained from $V^{k}$ is not affected by the rank-1 deflation. Briefly speaking, this is because adding a constant to $V^{k}$ through $\mathbf{1}$ has no effect on the $\operatorname*{arg\,max}$ computation used in the greedy policy. Indeed, the maximizer $\pi^{k+1}$ in equation 6 is the same as $\operatorname*{arg\,max}_{\pi}\{r^{\pi}+\gamma{\mathcal{P}}^{\pi}V^{k}\}$, produced by the (non-deflated) value iteration, when $W^{0}=V^{0}$. Having the term $v^{\top}\mathbf{1}W^{k}$ in the update $W^{k+1}=r^{\pi}+\gamma(\mathcal{P}^{\pi^{k+1}}-v^{\top}\mathbf{1})W^{k}$ adds the same constant to all states, so it does not change the maximizer of the next policy.
4.2 DDVI with Automatic Power Iteration
According to Fact 2, we can construct the deflation matrix $E_{s}$ for ${\mathcal{P}}^{\pi}$ if we have its top $s$ right eigenvectors. The top eigenvector $u_{1}=\mathbf{1}$ is known, which as we saw in Section 4.1, gives $E_{1}$ and rank-1 DDVI. We show that we can start with $s=1$ and run rank-1 DDVI, and using the calculations already done in DDVI updates, estimate the next top eigenvectors and increase the deflation rank.
@@ 4.2 DDVI with Automatic Power Iteration
Assume that the top $s+1$ eigenvalues of $\mathcal{P}^{\pi}$ have distinct magnitude, i.e., $1=\lambda_{1}>|\lambda_{2}|>\cdots>|\lambda_{s+1}|$. Let $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ be a rank-$s$ deflation matrix of $\mathcal{P}^{\pi}$, as in Fact 2. Consider the DDVI algorithm in equation 5 with $\alpha=1$ as in Corollary 3.2. If $V^{0}=0$, since
$\left(\mathcal{P}^{\pi}-E_{s}\right)(I-\gamma E_{s})^{-1}=\mathcal{P}^{\pi}-E_{s}$
(verified as equation 10 in Appendix B), we have
$W^{k}=\sum_{i=0}^{k-1}\gamma^{i}\left(\mathcal{P}^{\pi}-E_{s}\right)^{i}r^{\pi}.$
@@
Therefore, $\frac{1}{\gamma^{k}}(W^{k+1}-W^{k})=\left(\mathcal{P}^{\pi}-E_{s}\right)^{k}r^{\pi}$ is the iterates of a power iteration for the matrix $(\mathcal{P}^{\pi}-E_{s})$ starting from initial vector $r^{\pi}$. As discussed in Fact 2, the top eigenvalue of $(\mathcal{P}^{\pi}-E_{s})$ is $\lambda_{s+1}$. For large $k$, we expect $\frac{W^{k+1}-W^{k}}{\|W^{k+1}-W^{k}\|}\approx w$, where $w$ is the top eigenvector of $(\mathcal{P}^{\pi}-E_{s})$. With $w$, we can recover the $(s+1)$-th right eigenvector of $\mathcal{P}^{\pi}$ through the formula (Bru et al., 2012, Proposition 5)
$u_{s+1}=w-\sum^{s}_{i=1}\frac{\lambda_{i}v_{i}^{\top}w}{\lambda_{i}-\lambda_{s+1}}u_{i}.$
Algorithm 1 DDVI with AutoPI
1:Initialize $C,\epsilon$triangleright For example, $C=10$, $\epsilon=10^{-4}$
2:function DDVI(($s,V,K,\{\lambda_{i}\}^{s}_{i=1},\{u_{i}\}^{s}_{i=1}$))
3:     $V^{0}=V$, $c=0$, $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$ as in Fact 2
4:     for $k=0,\dots,K-1$ do
5:         if $c\geq C\,\,\text{and}\left|\frac{w^{k+1}}{\|w^{k+1}\|_{2}}-\frac{w^{k}}{\|w^{k}\|_{2}}\right|<\epsilon$ then
6:              $\lambda_{s+1}=(w^{k})^{\top}w^{k+1}/\left(\gamma\|w^{k}\|^{2}_{2}\right)$, $u_{s+1}=w^{k+1}-\sum^{s}_{i=1}\frac{\lambda_{i}v_{i}^{\top}w^{k+1}}{\lambda_{i}-\lambda_{s+1}}u_{i}$
7:              Return DDVI($s\!+\!1,V^{k}\!\!,K\!-\!c,\{\lambda_{i}\}^{s\!+\!1}_{i=1},\!\{u_{i}\}^{s\!+\!1}_{i=1}$)
8:         else
9:              $W^{k+1}=\gamma(\mathcal{P}^{\pi}-E_{s})V^{k}+r^{\pi}$, $V^{k+1}=(I-\gamma E_{s})^{-1}W^{k+1}$
10:              $w^{k+1}=W^{k+1}-W^{k}$, $w^{k}=W^{k}-W^{k-1}$
11:              $c=c+1$
12:     Return $V^{K}$
13:Initialize $V^{0}$ and $u_{1}=\mathbf{1}$, $\lambda_{1}=1$
14:DDVI$(1,V^{0},K,\lambda_{1},u_{1})$
@@
Leveraging this observation, DDVI with Automatic Power Iteration (AutoPI) computes an approximate rank-$s$ deflation matrix $E_{s}$ while performing DDVI: Start with a rank-$1$ deflation matrix, using the first right eigenvector $u_{1}=\mathbf{1}$, and carry out DDVI iterations. If a certain error criteria is satisfied, use $W^{k+1}-W^{k}$ to approximate the second right eigenvector $u_{2}$. Then, update the deflation matrix to be rank $2$, and gradually increase the deflation rank. We formalize this approach in Algorithm 1.
4.3 Rank-$s$ DDVI with the QR Iteration.
Algorithm 2 Rank-$s$ DDVI with QR Iteration
Initialize $\alpha$, $s$, and $V^{0}$
$\{\lambda_{i}\}^{s}_{i=1},\{u_{s}\}^{s}_{i=1}$ = QRiteration($\mathcal{P}^{\pi}$, s)
for $k=0,\dots,K-1$ do
The QR iteration of Algorithm 2 can be carried out in an automated manner similar to AutoPI. We refer to this as AutoQR and formally describe the algorithm in Appendix C.
5 Deflated Dynamics Temporal Difference Learning
@@
Many practical RL algorithms such as Temporal Difference (TD) Learning (Sutton, 1988; Tsitsiklis & V. Roy, 1997), Q-Learning (Watkins, 1989), Fitted Value Iteration (Gordon, 1995), and DQN (Mnih et al., 2015) can be viewed as sample-based variants of VI. Consequently, the slow convergence in the case of $\gamma\approx 1$ has also been observed for these algorithms by Szepesvári (1997); Even-Dar & Mansour (2003); Wainwright (2019) for TD Learning and by Munos & Szepesvári (2008); Farahmand et al. (2010); Chen & Jiang (2019); Fan et al. (2020) for Fitted VI and DQN. Here we introduce Deflated Dynamics Temporal Difference Learning (DDTD) as a sample-based variant of DDVI. To start, recall that the tabular temporal difference (TD) learning performs the updates
$V^{k+1}(X_{k})=V^{k}(X_{k})+\eta_{k}(X_{k})[r^{\pi}(X_{k})+\gamma V^{k}(X^{\prime}_{k})-V^{k}(X_{k})],$
where $\eta_{k}(X_{k})$ is a state-dependent stepsize, $(X_{k},X^{\prime}_{k})$ are random samples from the environment such that $X_{k}^{\prime}$ is the subsequent state following $X_{k}$, and $r^{\pi}(X_{k})$ is the expected immediate reward obtained by following policy $\pi$ from state $X_{k}$. In practice, we may choose a state-independent $\eta_{k}(X_{k})=\eta_{k}$, which would still result in convergence to the true value function if each state is visited often enough.
@@
TD is a sample-based variant of VI for PE. Two key ingredients of TD learning are the random coordinate updates and the temporal difference error $r^{\pi}(X)+\gamma V(X^{\prime})-V(X)$, whose conditional expectation is equal to $(T^{\pi}V-V)(X)$. Applying the same ingredients to DDVI, we obtain DDTD. The improved convergence rate for DDVI compared to VI suggests that DDTD may also exhibit improved convergence rates.
@@
$W^{k+1}(X_{k})$ $=W^{k}(X_{k})+\eta_{k}(X_{k})\big{[}(1-\alpha)V^{k}(X_{k})+\alpha\left(r^{\pi}(X_{k})+\gamma V(X^{\prime}_{k})-\gamma(E_{s}V^{k})(X_{k})\right)-W^{k}(X_{k})\big{]}$
@@
Let $\eta_{k}(x)=(\sum^{k}_{i=0}\mathbf{1}_{X_{i}=x})^{-1}$ when $X_{k}=x$. For $\alpha=1$, DDTD equation 8 converges to $V^{\pi}$ almost surely.
@@
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. Let $\eta_{k}(x)=C/(k+1)$ with the constant $C>\frac{n}{2\lambda_{\text{DDTD}}}$, where $\lambda_{\text{DDTD}}=\min_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\text{Re}(1-\gamma\lambda)$. For $\alpha=1$, DDTD equation 8 exhibits the rate $\mathbf{E}[\|V^{k}-V^{\pi}\|^{2}]=O(k^{-1})$.
@@
5.1 Implementation of DDTD
To implement DDTD practically, we take a hybrid of model-free and model-based approach for obtaining $E_{s}$. At each iteration, the agent uses the new samples to update an approximate model $\hat{\mathcal{P}}^{\pi}$ of the true dynamics. This updated approximate model is used to compute $E_{s}$. We update $E_{s}$ every $K$-th iterations since the change in $\hat{\mathcal{P}}^{\pi}$ and consequently the resulting $E_{s}$ at each iteration is small. Whenever a new $E_{s}$ is computed, we set $W$ to be $(I-\alpha\gamma E_{s})V$. This step ensures that $V$ smoothly converges to $V^{\pi}$ by updating all coordinates of $W$ at once based on new $E_{s}$. Also, we use the random sample reward $R$ in place of the expected reward $r^{\pi}(X_{k})$. We formally state the algorithm as Algorithm 3.